# Closure for Microscale Fluctuations in Chaotic Mixing Fronts

### v1

@@ III.6 Evolution of a mixing front
Now, we test the capacity of the proposed closure Eq. (35) to capture the evolution of a unidirectional mixing front, agitated by a chaotic microscale flow (the sine flow), and that spreads to large scales due to dispersion. Since $\bar{c}$ follows the macrodispersion equation with $\overline{\bm{u}}=0$, its solution for an initial sharp step function is
@@
The evolution of the local mean concentration is well captured by the dispersive equation (Fig. 7a)
@@
Eq. (52) accurately capture the evolution of scalar variance along the front at late time (Fig. 7b). At early times ($t\leq 5$), the prediction underestimates the spatial extent of the variance. This is probably due to the non-negligible transport of scalar variance at early time (the left hand side of Eq. (16)), which redistribute fluctuations.
@@ IV Conclusion
We investigated the statistical behaviour of mixing fronts in smooth chaotic flows in the presence of large-scale scalar gradients.
@@
We use the spectral description of mixing in time-uncorrelated flows of Kraichnan [1] to propose a new closure model relating the stretching statistics (mean and variance of the stretching rates) to the local variability of solute concentrations (the scalar variance), depending on the strength of large-scale scalar gradient and the Péclet number. The validity of the approach relies on three key properties of the flow. The first is the smoothness of the flow field, which ensures that stretching is a continuous, scale-independent process below the velocity-length scale. The second one is the scale separation induced by chaotic advection, which implies that the dissipation and production scales of the scalar variance are well separated, the Batchelor scale being well below the characteristic velocity length scale. The third is the rapid time-decorrelation of flow, which ensures that the evolution of stretching statistics can be envisioned as a Fokker-Planck equation (43). We show that this property is, however, not critical, and that a finite amount of flow persistence does not strongly affect the accuracy of the model. In addition, recent results [40] show that $\rho$ follows an Orstein-Uhlenbeck process for steady and transient flows, reducing to Eq. (22) in the Fickian case.
The variance of scalar fluctuations created by a chaotic flow in a mean scalar gradient is a universal function of a Péclet number (Fig. 6), defined as $Pe=\sigma^{2}_{\gamma}s_{i}^{2}/\kappa$, where $\sigma^{2}_{\gamma}$ is the variance of the stretching rates in the chaotic flow (typically of the same order of magnitude of the Lyapunov exponent $\gamma$), $s_{i}$ is a characteristic scale at which scalar variance is produced (typically of the order of the velocity length scale) and $\kappa$ the molecular diffusivity.
@@
We show that the proposed closure captures direct numerical simulations of advection-diffusion transport in a prototype chaotic flow, the sine flow, and is able to predict the evolution of mean and fluctuation of a solute mixing front. Numerical observation also suggests that local concentration statistics are essentially Gaussian, such that the modelling of the mean and variance of concentrations is sufficient to describe the behaviour of scalar plumes in smooth chaotic flows.
@@
The microscopic closure (35) is expected to apply in any smooth single-scale chaotic flow of low persistence. This is the case for steady laminar flows, for instance in Stokes flow through 3D porous media, where the velocity fluctuations induce chaotic advection in the 2D plane transverse to mean flow direction [41]. In such flows, fluid deformations occur on a time scale that is related to the velocity length scale and magnitude ($T\sim s_{v}/U$), such that Eq. (24) simplifies to $\gamma\sim U/s_{v}$. Thus, flow persistence is of order $Ku\sim 1$. In fact, the microscale $s_{m}$ has already been studied in the context of 3D heterogeneous flows at Darcy scale  Kapoor and Gelhar [42]. However, above pore-scale, advection is in general not chaotic [43], and $s_{m}$ is a non-trivial function of time and space, which strongly dependent on initial conditions. In contrast, below the Darcy scale, advection is chaotic and the microscale is a constant, independent of the solute plume geometry. Thus, the closure (35) is expected to be relevant in the context of up-scaling transport from pore to Darcy scale in porous media.
@@
More work is still required to estimate the value of $s_{i}$ in flows other than the sine flow and to confirm its physical origin and universality (Eq. (36)). Numerical test of the closure in 3D-time dependent chaotic flows is also of direct interest. Because it describes the amount of mixing at microscale, the proposed closure opens a new avenue to model non-conservative transport processes, such as reactive processes. Extension to bimolecular reactive transport system can be obtained by focusing on the concentration statistics of two reactive species, with fluctuation product $\overline{c_{A}^{\prime}c_{B}^{\prime}}$ rather than a single one $\overline{c^{\prime}c^{\prime}}$ as done here [44]. This will be the object of a future study.

### v2

@@ III.5 Evolution of a mixing front
We now test the capacity of the proposed closure Eq. (21) to capture the evolution in space and time of the scalar variance in a unidirectional mixing front, agitated by a chaotic microscale flow (the sine flow), and that spreads to large scales due to dispersion. Since $\bar{c}$ follows the macroscopci dispersion equation with $\overline{\bm{u}}=0$, its solution for an initial sharp step function is
@@
The evolution of the local mean concentration is well captured by the dispersive equation (Fig. 8a). The mean concentration is obtained numerically as $\bar{c}(x,t)=\langle c(x,y,t)\rangle_{y}$, where the average is taken on the direction $y$ and on a ensemble of 100 realizations of the random phases $(\phi_{i},\phi_{j})$.
@@
Numerically, we estimate the variance of scalar fluctuations as $\sigma^{2}_{c}(x,t)=\langle c(x,y,t)^{2}\rangle_{y}-\langle c(x,y,t)\rangle_{y}^{2}$. Eq. (41) accurately capture the evolution of scalar variance along the front at late time (Fig. 8b). At early times ($t\leq 5$), the prediction underestimates the spatial extent of the variance. This is probably due to the non-negligible transport of scalar variance at early time (the left hand side of Eq. (8)), which redistribute fluctuations.
@@ IV Conclusion
In this study, we have investigated the statistical behaviour of mixing fronts in chaotic flows, with the objective of describing and predicting the interactions between macroscopic scalar gradients driven by dispersion and microscale mixing controlled by stretching and folding. The statistics of concentration fluctuations below the characteristic velocity scale are described the spectral theory of mixing in time-uncorrelated flows of Kraichnan [24] while the macroscopic concentration gradients are described by Fickian dispersion. To relate these two descriptions, we identity an injection scale $s_{i}$ defined as the scale at which the characteristic time to destroy variance by dispersion is comparable to the characteristic time to destroy variance by stretching-enhanced mixing. This leads to a prediction of the concentration statistics across mixing fronts in chaotic flows. The theory captures accurately the results of numerical simulations of mixing fronts in periodic sine flows with no fitting parameters.
@@
The validity of the approach relies on three key properties of the flow. The first is the smoothness of the flow field, which ensures that stretching is a continuous, scale-independent process below the velocity-length scale and that dispersion is Fickian above the velocity scale. The second one is the scale separation induced by chaotic advection, which implies that the dissipation and production scales of the scalar variance are well separated, the Batchelor scale being well below the characteristic velocity length scale. The third is the rapid time-decorrelation of flow, which ensures that the evolution of stretching statistics can be envisioned as a Fokker-Planck equation (31). We show that this property is, however, not critical, and that a finite amount of flow persistence does not strongly affect the accuracy of the model. In addition, recent results [40] show that $\rho$ follows an Ornstein-Uhlenbeck process for steady and transient flows. Analytical results may thus also be within reach for flows with finite persistence.
@@
One key result of our analysis is that the variance of scalar fluctuations in a freely dispersing mixing front agitated by a smooth chaotic flow is a unique function of a Péclet number comparing dispersive to diffusive fluxes (Eq. (24)). An explicit expression for the closure is given by Eq. (21), and we show that it captures direct numerical simulations of advection-diffusion transport in a prototype chaotic flow, the sine flow, and is able to predict the evolution of mean and fluctuation of a solute mixing front. Numerical observation also suggests that local concentration statistics are essentially Gaussian, such that the modelling of the mean and variance of concentrations is sufficient to describe the behaviour of scalar plumes in smooth chaotic flows.
@@
The closure (21) is expected to apply in any smooth single-scale chaotic flow of moderate persistence. This is the case for steady laminar flows, for instance in Stokes flow through 3D porous media, where the velocity fluctuations induce chaotic advection in the 2D plane transverse to mean flow direction [41]. In such flows, fluid deformations occur on a time scale that is related to the velocity length scale and magnitude ($T\sim s_{v}/U$), such that Eq. (9) simplifies to $\gamma\sim U/s_{v}$. Thus, flow persistence is of order $Ku\sim 1$.
@@
Because it describes the amount of mixing at microscale in a macroscopically dispersing front, the proposed closure opens a new avenue to upscale non-conservative transport processes, such as reactive processes. Extension to bimolecular reactive transport system can be obtained by focusing on the concentration statistics of two reactive species, with fluctuation product $\overline{c_{A}^{\prime}c_{B}^{\prime}}$ rather than a single one $\overline{c^{\prime}c^{\prime}}$ as done here [42]. This will be the focus of a future study.
Additional work is also required to validate the proposed expression for $s_{i}$ (Eq. (19)) on other flows than the sine flow and to confirm its physical origin and significance in other context. Numerical test of the closure in 3D-time dependent chaotic flows is also of direct interest and could be investigated under the same framework.